# WavePulse: Real-time Content Analytics of Radio Livestreams

### v1

@@ Abstract.
Web content analytics, Radio livestreams, Large language models
copyright: noneconference: ; Preprint; Under Reviewccs: General and reference Measurementccs: Information systems Multimedia streamingccs: Computing methodologies Information extraction
@@ 1. Introduction
The rest of this paper is organized as follows. We first describe the WavePulse framework, followed by details about the data acquisition process. Next, we provide our qualitative and quantitative findings from the three case studies described above. We provide a discussion of related work, and conclude with potential directions of future research.
@@ 3.1. Case Study: Spread of a Political Narrative
Context and Background. A major focus area of the democracy group at this non-profit is to perform election monitoring worldwide. A specific current goal of theirs is to track election related claims across the U.S. Prior to WavePulse, their reach was limited in scope to social media posts, blogs, news articles, podcasts and comments.
@@
Manual Methodology: Using the above keywords, we first filtered transcripts of stations in Georgia. E.g., a filtered transcript should have "logic" AND "accuracy" AND "test" keywords. However, this resulted in only a few matches. We expanded our search nation-wide, which revealed 120 samples. As keywords acted as a broad net and we needed exact citations, we did an internal human study amongst four individuals. We split the 120 samples into four overlapping sets so that each transcript gets seen atleast twice. Then, we aggregated the responses and received a ground truth with 53 transcripts – 34 supporting the claim, 16 neutrally reporting and 3 actively debunking this claim. An excerpt from a June 2024 transcript that matched:
A matching transcript from Ohio (US) reads: “You know, here, the must-read of the day, [Claimant’s name] on how Georgia was stolen. A longtime spokesperson for President Trump, published a must-read analysis of the 2020 stolen Georgia election on [website], Wednesday. This is the most comprehensive investigative piece of the fraudulent election in Georgia to date. Biden allegedly received the most votes of a president in U.S. history, yet simultaneously could only sneak up in Georgia by some 13,558 votes. It didn’t add up. It just didn’t add up, and neither did the numbers in Fulton County.”
@@ 3.2. Case Study: Content Syndication Across Radio Stations
    In the remaining subgroups, extracted only the station names, such as KLMN and KOPQ, for each unique station.
@@
    Created bidirectional edges between stations in each subgroup (e.g., for stations KLMN, KOPQ, and KRST, edges were created between all pairs).
@@
To validate our results, we employed the Louvain Community Detection Algorithm (louvain_community_detect, ) on the US radio station graph, basing our analysis exclusively on edges derived from spoken content similarity. By coloring the resulting community detection graph according to station formats, we observed distinct clustering patterns where similar content types naturally grouped together while differing content types displayed greater separation in the network structure (See Fig. 13). These validation efforts substantially reinforced our confidence in WavePulse’s capabilities to accurately identify both narrative spread and content syndication patterns throughout radio networks.
@@ 3.3. Case Study: Presidential Candidates’ Favorability Trends
We isolated relevant text segments by keyword matching, carefully excluding instances of multiple candidate mentions to ensure sentiment clarity [^3] Name variations (e.g., ”Kamala” for Harris) were aggregated under primary identifiers for consistency.. The Twitter-roBERTa-base model (camacho-2022-tweetnlp, ), denoted as $\mathcal{S}$, served as the foundation for sentiment analysis, generating positive ($S_{\oplus}$), neutral ($S_{\odot}$), and negative ($S_{\ominus}$) sentiment counts. To distill these multifaceted sentiment counts into a single, comprehensible metric, we developed a normalized sentiment score $\bar{S}\in[0,1]$:
@@ 5. Discussion and Conclusion
The dataset may include PII about advertisers (e.g., names and contact information of organizational representatives) and show hosts. Given our focus on political topics, we have strived to avoid political bias by using politically neutral prompts and presenting case studies without subjective labeling. We acknowledge that the usage of LLMs in our analysis may inherently exhibit some biases due to their respective training data. We leave the study of any political bias to political scientists.
Limitations and Future Directions. Our analysis does not incorporate population data along with reach of each stations waves’ to calculate exposure to each station. Nielsen Audio sells exposure ratings and FCC hosts ground conductivity data. We also included only stations which are livestreamed over the Internet. Terrestrial-only radio broadcasts would require dedicated hardware (an antenna, transceiver, and recording equipment). Finally, while WavePulse is widely applicable, our analysis derives results and conclusions from only US radios. An important direction of future work is to broaden the scope to worldwide radio livestreams; due to the multilingual nature of LLMs we anticipate our system to scale up with no significant design changes.
While we are confident about WavePulse’s scalability and error rates, our analyses should be considered within the context of case studies. A comprehensive evaluation would require interdisciplinary collaboration and usage-based assessment. Our case studies demonstrate a non-exhaustive variants of general-purpose tasks – search, finding syndicate network, sentiment analysis – that WavePulse can perform.
@@ A.1. Data Collection Pipeline
A.1.1. Radio Streamer
This component is used to stream and record audio streams from multiple parallel radio broadcasts. It takes as input a configurable schedule in the form of a JSON file, for when to stream and from which radio. There will be one entry per radio stream, its live URL, name, record times, and the state in which it is located along with a list start and end times for streaming for that particular radio stream, as shown in Listing 1. The component processes the schedule to get unique start times across all the stations and durations specific for each station and then create cron jobs appropriately to achieve its objective.
@@ A.1.1. Radio Streamer
A.1.2. Audio Processor
This component is responsible for the transcription, adding punctuation and capitalization to the text, providing time stamps, and diarization of the speakers of the recorded audio files. We tried multiple ASR models for transcription like facebook’s MMS-1B, Nvidia’s Parakeet-RNNT-1.1B and OpenAI’s whisper-large-v3. Fig. 2 has sample outputs of these models for same input audio. We found word error rate of mms-1b to be significantly higher compared to Parakeet-RNNT and whisper-large-v3, for the latter two it was comparable. Ultimately we decided to go ahead with WhisperX implementation of Whisper as it provides a built-in pipeline for transcription using Whisper, accurate timestamps using Wav2Vec2 and Speaker Diarization using PyAnnote at a reasonable inference speed. Listing 1. shows a snippet of output using WhisperX pipeline.
@@ A.1.2. Audio Processor
WavePulse Deployment
Radio-streamer Configuration:
Streams $:396\text{ on }64\text{ cores}$
@@
[…]
Resource Utilization:
[…]
Minimal Deployment Requirements (100 streams):
Storage $:200\text{ GB}$
Computing $:2\text{ GPUs }+24\text{ CPU cores}$
(16 streaming, 4 per GPU)
Figure 9. Line chart showing the addition of radio stations over time.
@@ A.3. Dense Embedding for Summaries
A.3.1. Overview
Once the conversation segments are generated and summarized, the next step is to create embeddings for each summary. These embeddings serve as high-dimensional vector representations that capture the semantic meaning of the text. The embeddings are generated using BGE-M3 (bge-m3, ). By converting textual summaries into vectors, we enable efficient operations like similarity comparisons, clustering, and semantic search.
A.3.2. Mathematical Representation
The embedding function, denoted by $f$, transforms a given summary $S$ into a high-dimensional vector $\mathbf{v}\in\mathbb{R}^{d}$, where $d$ represents the number of dimensions in the vector space. The process can be mathematically described as:
@@ A.3.2. Mathematical Representation
A.3.3. Cosine Similarity of Embeddings
To compare the semantic similarity between two summary embeddings, cosine similarity is used. The cosine similarity between two vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, representing summaries $S_{i}$ and $S_{j}$, is given by:
@@ A.3.3. Cosine Similarity of Embeddings
A.3.4. Asynchronous Embedding Generation
Given the large number of summaries, the embeddings are generated asynchronously to improve performance and scalability. By parallelizing the generation process, we can significantly reduce the processing time, ensuring that embeddings are computed efficiently for each summary.
@@ A.3.4. Asynchronous Embedding Generation
A.3.5. Usage of Embeddings
The generated embeddings are then used for multiple downstream tasks, such as:
@@ A.4. Semantic Similarity Search with FAISS
A.4.1. Why FAISS?
FAISS (Facebook AI Similarity Search) was chosen for its ability to perform efficient nearest-neighbor search in high-dimensional vector spaces. Given the large number of summaries and their associated embeddings, FAISS provides a highly scalable solution for searching through these embeddings.
A.4.2. FAISS Workflow
The FAISS workflow consists of the following steps:
@@ A.5. Finding Unique Narrative Network
A.6.1. Home Page
The Home Page functions as the central navigation hub for the Wave-Pulse.io application, providing users with access to various features and data visualization options.
A.6.2. Map UI
The Map UI is built upon the ComposableMap component from ’react-simple-maps’, offering users an interactive exploration of the United States map. This visualization includes state boundaries, county outlines, and markers representing population centers and radio station coverage areas.
@@ A.6.2. Map UI
A.6.3. Plots UI
The Plots UI component features two primary line charts: a nationwide combined sentiment analysis and a nationwide sentiment count. These charts are designed to adjust based on user interactions, providing a responsive data exploration experience. To add depth to the visualizations, entropy is incorporated, providing context to the plotted data and illustrating the degree of uncertainty in the sentiment analysis.
@@ A.6.3. Plots UI
Figure 16. Comprehensive sentiment plot for all parties and candidates, covering the period from June 26, 2024, to August 7, 2024.
[…]
a figure
@@
A.7.1. GitHub Pages Frontend
GitHub Pages serves as the hosting platform for the frontend, providing key advantages:
Performance
GitHub Pages delivers pre-built assets directly to the browser, resulting in fast load times for users. This approach eliminates the need for server-side rendering, enhancing the responsiveness of the application.
Automated Deployment
Updates to the frontend trigger automatic deployment processes using GitHub Actions. This automation streamlines the development workflow and ensures that new features and improvements are made available to users.
Advantages
  • Cost-effectiveness: GitHub Pages offers free hosting for public repositories, reducing operational costs.
  • Seamless integration: The tight integration with the GitHub ecosystem facilitates a smooth development and deployment process.
  • Version control: Inherent version control capabilities allow for tracking of changes and rollbacks if necessary.
A.7.2. DigitalOcean Backend
The DigitalOcean App Platform is employed to handle the Django backend and PostgreSQL database, offering a robust and scalable solution for server-side operations.
Functionality
@@ A.7.2. DigitalOcean Backend
A.8. Extension: Scraping Fact Checks
The fact-checking program is a system designed for the automated collection, processing, and post-processing of fact-checking articles from various reputable websites. Developed using Python and leveraging the Scrapy framework for web scraping, the system incorporates specialized spiders for specific fact-checking websites, including FactCheck.org, Lead Stories, Politifact, Snopes, and TruthOrFiction. Although we do not use it in WavePulse, matching fact checks from authoritative resources with transcript embeddings can help automate fact-checking.
A.8.1. Websites Overview
Our team developed a Fact Check Web Crawler capable of scraping fact-checking articles from multiple authoritative websites. The data collection period spans from January 1, 2020, to August 6, 2024, providing a substantial dataset for analysis. To ensure ethical and legal compliance, we reviewed the terms of service for each website prior to data collection.
A.8.2. Scrapy Implementation
The core of our fact-checking system utilizes Scrapy spiders to scrape fact-checking articles. These spiders are designed with flexibility, supporting various filtering options including date range, keywords, tags, and pagination. This adaptability allows for targeted data collection based on specific research needs.
@@ A.8.2. Scrapy Implementation
A.8.3. Deduplication Process
To ensure data integrity and prevent redundancy, we implemented a deduplication module. This component is designed to identify and manage duplicate fact-checking articles by conducting analysis of textual content and publication dates.
@@ A.8.3. Deduplication Process
Figure 19. Illustrative example of misinformation fact-checked across multiple reputable websites, demonstrating the need for effective deduplication.
[…]
a figure
A.8.4. Results Visualization
To facilitate understanding and analysis of the collected fact-checking data, we developed visualization tools, including word clouds and histograms.
Histogram Analysis
@@ A.8.4. Results Visualization
Figure 20. Comprehensive histogram illustrating the frequency of occurrence for all fact-checking "rulings" across the scraped websites for the year 2024.
[…]
a figure
@@
[…]
a figure
Figure 23. Word cloud showing frequency of occurrence of significant words with the ruling "Mostly True", "Mixed", or "Mostly False" across all the data scraped from fact-checking websites for 2024.
[…]
a figure
Figure 24. Word cloud showing frequency of occurrence of significant words with the remaining rulings such as "Satire" or "Outdated" across all the data scraped from fact-checking websites for 2024.
@@ A.9. Extension: Audio Classifier
KFXZ Lafayette, LA KSYL Alexandria, LA KWLA Anacoco, LA

### v2

@@ Abstract.
Web content analytics; Radio livestreams; Large language models
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714810isbn: 979-8-4007-1274-6/25/04ccs: General and reference Measurementccs: Information systems Multimedia streamingccs: Computing methodologies Information extraction
@@ 1. Introduction
The rest of this paper is organized as follows. We first describe the WavePulse framework, followed by details about the data acquisition process. Next, we provide our qualitative and quantitative findings from the three case studies described above. We provide a discussion of related work, and conclude with future directions.
@@ 3.1. Case Study: Spread of a Political Narrative
Context and Background. A major focus area of the democracy group is to perform election monitoring worldwide. A specific current goal of theirs is to track election related claims across the U.S. Prior to WavePulse, their reach was limited in scope to social media posts, blogs, news articles, and podcasts.
@@
Manual Methodology: Using these keywords, we first filtered transcripts of stations in only Georgia. E.g., a filtered transcript should have "logic" AND "accuracy" AND "test" keywords. However, this resulted in only a few matches. We expanded our search nation-wide, which revealed 120 samples. As keywords acted as a broad net and we needed exact citations, we did an internal human study amongst four individuals. We split the 120 samples into four overlapping sets so that each transcript verified twice. Then, we aggregated the responses into 53 ground truth transcripts: 34 promoting the claim, 16 neutrally reporting and 3 actively debunking.
A matched June 2024 transcript from Ohio (US) read: “You know, here, the must-read of the day, [Claimant’s name] on how Georgia was stolen. A longtime spokesperson for President Trump, published a must-read analysis of the 2020 stolen Georgia election on [website], Wednesday. This is the most comprehensive investigative piece of the fraudulent election in Georgia to date. Biden allegedly received the most votes of a president in U.S. history, yet simultaneously could only sneak up in Georgia by some 13,558 votes. It didn’t add up. It just didn’t add up, and neither did the numbers in Fulton County.”
@@ 3.2. Case Study: Content Syndication Across Radio Stations
    In the remaining subgroups, extracted only the station names, such as KLMN, for each unique station.
@@
    Created bidirectional edges between stations in each subgroup.
@@
To validate our results, we employed the Louvain Community Detection Algorithm (louvain_community_detect, ) on the US radio station graph, basing our analysis exclusively on edges derived from spoken content similarity. By coloring the resulting community detection graph according to station formats, we observed distinct clustering patterns where similar content types naturally grouped together while differing content types displayed greater separation in the network structure (See Fig. 12). This validation reinforced WavePulse’s capabilities to accurately identify both narrative spread and content syndication patterns throughout radio networks.
@@ 3.3. Case Study: Presidential Candidates’ Favorability Trends
We isolated relevant text segments by keyword matching, carefully excluding instances of multiple candidate mentions to ensure sentiment clarity [^3] Name variations (e.g., ”Kamala” for Harris) were aggregated under primary identifiers.. The Twitter-roBERTa-base model (camacho-2022-tweetnlp, ), denoted as $\mathcal{S}$, served as the foundation for sentiment analysis, generating positive ($S_{\oplus}$), neutral ($S_{\odot}$), and negative ($S_{\ominus}$) sentiment counts. To distill these multifaceted sentiment counts into a single, comprehensible metric, we developed a normalized sentiment score $\bar{S}\in[0,1]$:
@@ 5. Discussion and Conclusion
The dataset may include info about advertisers (e.g., names and contact information of organizational representatives) and show hosts. Given our focus on political topics, we have strived to avoid political bias by using politically neutral prompts and presenting case studies without subjective labeling. We acknowledge that the usage of LLMs in our analysis may inherently exhibit some biases due to their respective training data. We leave the study of any political bias to political scientists.
Limitations and Future Directions. Our analysis does not incorporate population data along with reach of each stations waves’ to calculate exposure to each station. Nielsen Audio sells exposure ratings and FCC hosts ground conductivity data.
We also included only stations which are livestreamed over the Internet. Terrestrial-only radio broadcasts would require dedicated hardware (an antenna, transceiver, and recording equipment). Finally, while WavePulse is widely applicable, our analysis derives results and conclusions from only US radios. An important direction of future work is to broaden the scope to worldwide radio livestreams; due to the multilingual nature of LLMs we anticipate our system to scale up with no significant design changes.
@@
Acknowledgment
This work was supported in part by SATC grants #2154118 and #2154119 awarded by the National Science Foundation, and by a gift grant by the Google Cyber NYC Initiative.
@@ A.1. Data Collection Pipeline
Radio Streamer This component is used to stream and record audio streams from multiple parallel radio broadcasts. It takes as input a configurable schedule in the form of a JSON file, for when to stream and from which radio. There will be one entry per radio stream, its live URL, name, record times, and the state in which it is located along with a list start and end times for streaming for that particular radio stream, as shown in Listing 1. The component processes the schedule to get unique start times across all the stations and durations specific for each station and then create cron jobs appropriately to achieve its objective.
@@
Audio Processor This component is responsible for the transcription, adding punctuation and capitalization to the text, providing time stamps, and diarization of the speakers of the recorded audio files. We tried multiple ASR models for transcription like facebook’s MMS-1B, Nvidia’s Parakeet-RNNT-1.1B and OpenAI’s whisper-large-v3. Fig. 2 has sample outputs of these models for same input audio. We found word error rate of mms-1b to be significantly higher compared to Parakeet-RNNT and whisper-large-v3, for the latter two it was comparable. Ultimately we decided to go ahead with WhisperX implementation of Whisper as it provides a built-in pipeline for transcription using Whisper, accurate timestamps using Wav2Vec2 and Speaker Diarization using PyAnnote at a reasonable inference speed. Listing 1. shows a snippet of output using WhisperX pipeline.
@@
WavePulse Deployment.
@@
Figure 8. Line chart showing the addition of radio stations over time and plugged into WavePulse.
@@ A.3. Dense Embedding for Summaries
Overview. Once the conversation segments are generated and summarized, the next step is to create embeddings for each summary. These embeddings serve as high-dimensional vector representations that capture the semantic meaning of the text. The embeddings are generated using BGE-M3 (bge-m3, ). By converting textual summaries into vectors, we enable efficient operations like similarity comparisons, clustering, and semantic search.
Mathematical Representation. The embedding function, denoted by $f$, transforms a given summary $S$ into a high-dimensional vector $\mathbf{v}\in\mathbb{R}^{d}$, where $d$ represents the number of dimensions in the vector space. The process can be mathematically described as:
@@
Cosine Similarity of Embeddings. To compare the semantic similarity between two summary embeddings, cosine similarity is used. The cosine similarity between two vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, representing summaries $S_{i}$ and $S_{j}$, is given by:
@@
Asynchronous Embedding Generation. Given the large number of summaries, the embeddings are generated asynchronously to improve performance and scalability. By parallelizing the generation process, we can significantly reduce the processing time, ensuring that embeddings are computed efficiently for each summary.
@@
Usage of Embeddings. The generated embeddings are then used for multiple downstream tasks, such as:
@@ A.4. Semantic Similarity Search with FAISS
Why FAISS? FAISS (Facebook AI Similarity Search) was chosen for its ability to perform efficient nearest-neighbor search in high-dimensional vector spaces. Given the large number of summaries and their associated embeddings, FAISS provides a highly scalable solution for searching through these embeddings.
FAISS Workflow The FAISS workflow consists of the following steps:
@@ A.5. Finding Unique Narrative Network
Home Page. The Home Page functions as the central navigation hub for the Wave-Pulse.io application, providing users with access to various features and data visualization options.
Map UI. The Map UI is built upon the ComposableMap component from ’react-simple-maps’, offering users an interactive exploration of the United States map. This visualization includes state boundaries, county outlines, and markers representing population centers and radio station coverage areas.
@@
Plots UI. This component features two primary line charts: a nationwide combined sentiment analysis and a nationwide sentiment count. These charts are designed to adjust based on user interactions, providing a responsive data exploration experience. To add depth to the visualizations, entropy is incorporated, providing context to the plotted data and illustrating the degree of uncertainty in the sentiment analysis.
@@
DigitalOcean Backend. The DO App Platform is employed to handle the Django backend and PostgreSQL database, offering a robust and scalable solution for server-side operations.
Functionality.
@@
A.8. Extension: Scraping Fact Checks.
The fact-checking module is an extension designed for the automated collection, processing, and post-processing of fact-checking articles from various reputable websites. Developed using Python and leveraging the Scrapy framework for web scraping, the system incorporates specialized spiders for specific fact-checking websites, including FactCheck.org, Lead Stories, Politifact, Snopes, and TruthOrFiction. Although we do not use it in WavePulse, matching fact checks from authoritative resources with transcript embeddings can help automate fact-checking.
Websites Overview. Our team developed a Fact Check Web Crawler capable of scraping fact-checking articles from multiple authoritative websites. The data collection period spans from January 1, 2020, to August 6, 2024, providing a substantial dataset for analysis. To ensure ethical and legal compliance, we reviewed the terms of service for each website prior to data collection.
Scrapy Implementation. The core of our fact-checking system utilizes Scrapy spiders to scrape fact-checking articles. These spiders are designed with flexibility, supporting various filtering options including date range, keywords, tags, and pagination. This adaptability allows for targeted data collection based on specific research needs.
@@
Deduplication Process. To ensure data integrity and prevent redundancy, we implemented a deduplication module. This component is designed to identify and manage duplicate fact-checking articles by conducting analysis of textual content and publication dates.
@@
Results Visualization. To facilitate understanding and analysis of the collected fact-checking data, we developed visualization tools, including word clouds and histograms.
Histogram Analysis.
@@ A.9. Extension: Audio Classifier
KFXZ Lafayette, LA KSYL Alexandria, LA KWLA Anacoco, LA